# Inverse Problem Sampling in Latent Space Using Sequential Monte Carlo

### v1

@@ Appendix B Proposal Distribution Scaling Coefficients
Recall that our proposal distribution (Eq. 4 in the main text) is made of two elements. These elements are scaled by two coefficients, $\gamma_{t}$ and $\lambda_{t}$. Here, we provide an explicit formula for these coefficients. We found that our proposed scaling works better than common procedures used in the literature. For consistency with baseline methods, we also used our proposed scaling approach for DPS and TDS, since these methods apply a similar update rule. We tried to use it for PSLD and Resample but it didn’t work well for these baselines.
Let ${\mathbf{g}}_{t}^{1}\coloneqq\nabla_{{\mathbf{z}}_{t}}||{\mathbf{y}}_{0}-{\mathcal{A}}({\mathcal{D}}(\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t})))||_{2}^{2}$, and ${\mathbf{g}}_{t}^{2}\coloneqq\nabla_{\bm{\mathbf{\mu}}_{\theta}(\hat{{\mathbf{z}}}_{t},t)}||{\mathbf{y}}_{t}-{\mathcal{A}}({\mathcal{D}}(\bm{\mathbf{\mu}}_{\theta}(\hat{{\mathbf{z}}}_{t},t)))||_{2}^{2}$. We set the scaling coefficients $\gamma_{t}$ and $\lambda_{t}$ the same and according to the following scheme: $\gamma_{t}=\kappa_{1}\cdot(1-\bar{\alpha}_{t})^{\nu_{1}}\cdot\frac{1}{max(||{\mathbf{g}}_{t}^{1}||_{2}^{2},1)}$, and similarly $\lambda_{t}=\kappa_{2}\cdot(1-\bar{\alpha}_{t})^{\nu_{2}}\cdot\frac{1}{max(||{\mathbf{g}}_{t}^{2}||_{2}^{2},1)}$. Here $\{\nu_{1},\nu_{2}\}$ are hyperparameters that controls the effect of the variance scaling, and $\{\kappa_{1},\kappa_{2}\}$ scale the entire term. In practice, since we use either ${\mathbf{g}}_{t}^{1}$ or ${\mathbf{g}}_{t}^{2}$, but not both, we have only one set of hyperparameters which is used for both $\gamma_{t}$ and $\lambda_{t}$. That is, $\nu_{1}\coloneqq\nu_{2}$ and $\kappa_{1}\coloneqq\kappa_{2}$.
Appendix C Forward Process
In Figure 5 we present the evolution of the auxiliary labels ${\mathbf{y}}_{t}$ over time as part of the forward process according to our proposed sampling procedure in Section 4.2, steps (a) & (b). From the figure, we observe a gradual cleaning of noise in the auxiliary labels when advancing from time $t=999$ to time $t=0$.
Appendix D Full Results
Table 4: ImageNet. Box in-painting on $1024$ test examples.
@@ Appendix D Full Results
Latent DPS 65.45 7.918 19.19 0.623 0.407
Latent TDS 65.03 7.872 19.21 0.623 0.406
Resample 90.32 8.464 18.16 0.695 0.318
PSLD 79.90 9.268 17.48 0.583 0.410
LD-SMC (Ours) 51.81 5.103 18.87 0.599 0.355
Table 5: ImageNet. Gaussian debluring on $1024$ test examples.
@@
Latent DPS 52.48 6.855 23.61 0.615 0.383
Latent TDS 50.82 6.695 23.57 0.614 0.379
Resample 46.45 7.411 24.36 0.639 0.353
PSLD 79.31 7.972 21.38 0.483 0.474
LD-SMC (Ours) 52.17 6.789 23.60 0.614 0.382
Table 6: ImageNet. Super Resolution (8×8$\times$8 ×) on $1024$ test examples.
@@
Latent DPS 61.02 6.514 21.65 0.523 0.439
Latent TDS 58.73 7.157 21.45 0.515 0.454
Resample 87.65 8.290 22.05 0.532 0.491
PSLD 78.56 7.000 21.54 0.516 0.467
LD-SMC (Ours) 59.27 6.423 21.64 0.521 0.437
Table 7: FFHQ. Box in-painting on $1024$ test examples.
@@
Latent DPS 39.81 7.592 24.15 0.814 0.236
Latent TDS 39.57 7.602 24.24 0.814 0.236
Resample 86.79 7.142 19.75 0.815 0.230
PSLD 47.51 7.480 22.70 0.722 0.312
LD-SMC (Ours) 37.14 7.520 24.08 0.817 0.224
Table 8: FFHQ. Gaussian debluring on $1024$ test examples.
@@
Latent DPS 31.81 6.813 26.25 0.709 0.285
Latent TDS 33.19 6.879 26.13 0.705 0.288
@@
LD-SMC (Ours) 32.18 6.566 26.60 0.721 0.280
Table 9: FFHQ. Super resolution (8×8$\times$8 ×) on $1024$ test examples.
@@
LD-SMC (Ours) 30.37 6.456 25.42 0.698 0.284
Appendix E Image Reconstructions
Figure 6: Comparison between LD-SMC and baseline methods on inpainting of images from the FFHQ dataset.
Figure 7: Gaussian debluring. LD-SMC reconstruction of images from FFHQ (left) and ImageNet (right).
Figure 8: Super resolution. LD-SMC reconstruction of images from FFHQ (left) and ImageNet (right).

### v2

@@ Appendix D Proposal Distribution Scaling Coefficients
Recall that our proposal distribution (Eq. 7 in the main text) is made up of two elements. These elements are scaled by two coefficients, $\gamma_{t}$ and $\lambda_{t}$. Here, we provide an explicit formula for these coefficients. We found that our proposed scaling works better than common procedures used in the literature. For consistency with baseline methods, we also used our proposed scaling approach for DPS and TDS (which corresponds to setting $s=0$), since these methods apply a similar update rule. We tried to use it for other baselines but it did not work well for them.
Let ${\mathbf{g}}_{t}^{1}\coloneqq\nabla_{{\mathbf{z}}_{t}}||{\mathbf{y}}_{0}-{\mathcal{A}}({\mathcal{D}}(\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t})))||_{2}^{2}$, and ${\mathbf{g}}_{t}^{2}\coloneqq\nabla_{\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t+1},t+1)}||{\mathbf{y}}_{t}-{\mathcal{A}}({\mathcal{D}}(\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t+1},t+1)))||_{2}^{2}$. We set the scaling coefficients $\gamma_{t}$ and $\lambda_{t}$ according to the following scheme: $\gamma_{t}=\mathbbm{1}_{[t\geq s]}\cdot\kappa_{1}\cdot\frac{1}{max(||{\mathbf{g}}_{t}^{1}||_{2}^{2},1)}+\mathbbm{1}_{[t<s]}\cdot\kappa_{2}\cdot(1-\rho)\cdot\frac{1}{max(||{\mathbf{g}}_{t}^{1}||_{2}^{2},1)}$, and $\lambda_{t}=\mathbbm{1}_{[t<s]}\cdot\kappa_{2}\cdot\rho\cdot\frac{1}{max(||{\mathbf{g}}_{t}^{2}||_{2}^{2},1)}$, where $s,\rho,\kappa_{1},\kappa_{2}$ are all hyper-parameters.
Appendix E Computational Cost
We add here a comparison between the compared methods in average run-time (seconds) and memory (GB) over 10 trials for sampling a single image on ImageNet box inpainting task. For TDS and PFLD we use 5 particles. For LD-SMC, we inspect several variants that differ in the number of particles and Gibbs iterations. The results are shown in Table 4. From the table, the run-time is roughly linear in the number of particles and Gibbs iterations. Yet, importantly, it can be controlled by the practitioner to trade off performance (which can be good with one Gibbs iteration and one particle) and computational demand. In addition, we note that our code is not properly optimized and improvements can be made to it.
Table 4: Average run time and memory demand on ImageNet box inpainting task.
Method Run Time (Sec.) Memory (GB)
LDPS 105.5 8.123
LTDS 418.5 19.86
ReSample 333.4 5.769
PSLD 144.5 9.473
PFLD 469.8 26.51
LatentDAPS 67.82 4.745
LD-SMC (1 particle) 136.3 9.213
LD-SMC (3 particles) 375.1 15.11
LD-SMC (5 particles) 537.2 21.16
LD-SMC (10 particles) 1013. 35.78
LD-SMC (1 particle; 2 Gibbs iterations) 271.2 9.213
LD-SMC (1 particle; 4 Gibbs iterations) 541.0 9.213
Appendix F Further Ablation Studies
Number of Gibbs iterations. Recall that in the main text we used only one Gibbs step throughout. Here we examine the effect of using more Gibbs steps in Figure 6. We do so on two tasks, the first is super-resolution and the second in Gaussian deblurring, both on ImageNet. In super-resolution, we observe an improvement in the PSNR with the number of Gibbs steps while suffering from reduction in the FID. Interestingly, this experiment shows that we can obtain a comparable PSNR to ReSample while having a substantially better FID compared to it on this task. In Gaussian deblurring there is a less clear trend, nevertheless, the figure shows that the FID can be improved while maintaining roughly the same PSNR value. We speculate that the difference between these two cases stems from the different $s$ values used in these experiments. Taking $s>0$ tends to generate smoother images that are better aligned with higher values of distortion metrics such as PSNR.
Proposal distributions. In Table 5 we compare different choices of proposal distributions. Specifically, we compare to (1) the naive choice of using the prior as a proposal distribution, i.e., a bootstrap filter (Gordon et al., 1993); and (2) taking the DPS step as a proposal distribution, which corresponds to setting $s=0$. The table shows a clear advantage to the proposal distribution used in this study and presented in Section 4.2.2.
Table 5: FFHQ. Proposal distribution ablation. Box inpainting on $1024$ test examples using $N=5$ particles.
@@ Appendix F Further Ablation Studies
Prior proposal 62.35 7.628 12.60 0.365 0.671
DPS proposal 39.94 7.558 24.20 0.814 0.235
LD-SMC 33.87 7.066 24.10 0.821 0.211
Appendix G Forward Process
In Figure 7 we present the evolution of the auxiliary labels ${\mathbf{y}}_{t}$ over time as part of the forward process according to our proposed sampling procedure in Section 4.2, steps (a) & (b). From the figure, we observe a gradual cleaning of noise in the auxiliary labels when advancing from time $t=1000$ to time $t=0$.
Appendix H Full Results
Table 6: ImageNet. Box inpainting on $1024$ test examples.
@@ Appendix H Full Results
LDPS 65.04 7.935 19.51 0.665 0.379
LTDS 64.74 7.907 19.49 0.665 0.378
ReSample 90.32 8.464 18.16 0.695 0.318
PSLD 71.15 8.042 18.10 0.565 0.434
PFLD 72.83 7.933 18.09 0.564 0.436
LatentDAPS 98.24 11.36 19.98 0.704 0.394
LD-SMC (1 particle) 51.49 6.878 19.38 0.672 0.326
LD-SMC (5 particles) 50.67 6.891 19.42 0.672 0.325
Table 7: ImageNet. Free-form inpainting on $1024$ test examples.
@@
LDPS 53.47 7.867 23.76 0.722 0.334
LTDS 52.75 7.884 23.74 0.721 0.334
ReSample 44.15 7.104 22.78 0.778 0.248
PSLD 62.38 8.037 21.65 0.610 0.411
PFLD 63.34 8.026 21.59 0.609 0.414
LatentDAPS 68.65 9.625 24.75 0.766 0.330
LD-SMC (1 particle) 38.21 6.969 23.54 0.727 0.285
LD-SMC (5 particles) 36.18 6.671 23.43 0.727 0.278
Table 8: ImageNet. Gaussian deblurring on $1024$ test examples.
@@
LDPS 52.48 6.855 23.61 0.615 0.383
LTDS 50.82 6.695 23.57 0.614 0.379
ReSample 46.45 7.411 24.36 0.639 0.353
PSLD 60.68 6.599 21.94 0.506 0.417
PFLD 60.94 6.733 21.78 0.496 0.421
LatentDAPS 77.09 10.55 24.44 0.659 0.417
LD-SMC (1 particle) 52.48 6.855 23.61 0.615 0.383
LD-SMC (5 particles) 52.29 6.791 23.61 0.615 0.383
Table 9: ImageNet. Super-resolution (8×8$\times$8 ×) on $1024$ test examples.
@@
LDPS 61.02 6.514 21.65 0.523 0.439
LTDS 59.12 6.270 21.59 0.520 0.435
ReSample 87.65 8.290 22.05 0.532 0.491
PSLD 66.56 7.669 20.83 0.480 0.489
PFLD 64.72 7.685 20.83 0.479 0.492
LatentDAPS 104.6 12.68 22.38 0.566 0.489
LD-SMC (1 particle) 58.06 6.243 21.51 0.524 0.434
LD-SMC (5 particles) 57.89 6.238 21.50 0.520 0.433
Table 10: FFHQ. Box inpainting on $1024$ test examples.
Perceptual Quality Distortion
LDPS 39.81 7.592 24.15 0.814 0.236
LTDS 39.57 7.602 24.24 0.814 0.236
ReSample 86.79 7.142 19.75 0.815 0.230
PSLD 39.68 6.544 22.31 0.774 0.246
PFLD 39.06 6.509 22.40 0.774 0.245
LatentDAPS 60.24 9.999 24.91 0.838 0.257
LD-SMC (1 particle) 33.37 7.032 23.98 0.819 0.212
LD-SMC (5 particles) 33.87 7.066 24.10 0.821 0.211
Table 11: FFHQ. Free-form inpainting on $1024$ test examples.
Perceptual Quality Distortion
LDPS 40.17 7.609 27.95 0.858 0.212
LTDS 39.78 7.578 27.95 0.858 0.212
ReSample 37.01 6.622 26.31 0.891 0.151
PSLD 36.26 6.835 26.19 0.823 0.216
PFLD 36.43 6.817 26.35 0.825 0.215
LatentDAPS 54.40 8.766 29.23 0.883 0.223
LD-SMC (1 particle) 33.67 7.034 27.35 0.859 0.194
LD-SMC (5 particles) 33.60 7.021 27.35 0.859 0.194
Table 12: FFHQ. Gaussian deblurring on $1024$ test examples.
Perceptual Quality Distortion
LDPS 29.30 6.538 28.03 0.775 0.237
LTDS 30.23 6.553 27.93 0.772 0.238
@@
PFLD 37.16 6.751 23.96 0.628 0.343
LatentDAPS 54.28 9.496 29.70 0.831 0.283
LD-SMC (1 particle) 29.19 6.575 28.37 0.789 0.232
LD-SMC (5 particles) 29.47 6.528 28.34 0.787 0.233
Table 13: FFHQ. Super-resolution (8×8$\times$8 ×) on $1024$ test examples.
@@
PFLD 38.11 6.832 23.69 0.617 0.345
LatentDAPS 70.24 10.17 26.87 0.760 0.344
LD-SMC (1 particle) 30.02 6.426 25.52 0.706 0.277
LD-SMC (5 particles) 30.62 6.455 25.49 0.703 0.278
Appendix I Additional Image Reconstructions
Figure 8: Comparison between LD-SMC and baseline methods on inpainting of FFHQ images.
Figure 9: Gaussian deblurring. LD-SMC reconstruction of images from FFHQ (left) and ImageNet (right).
Figure 10: Super-resolution. LD-SMC reconstruction of images from FFHQ (left) and ImageNet (right).